# Multi-Agent Reinforcement Learning for Connected and Automated Vehicles Control: Recent Advancements and Future Prospects

### v1

@@ Abstract
Connected and automated vehicles (CAVs) have emerged as a potential solution to the future challenges of developing safe, efficient, and eco-friendly transportation systems. However, CAV control presents significant challenges, given the complexity of interconnectivity and coordination required among the vehicles. To address this, multi-agent reinforcement learning (MARL), with its notable advancements in addressing complex problems in autonomous driving, robotics, and human-vehicle interaction, has emerged as a promising tool for enhancing the capabilities of CAVs. However, there is a notable absence of current reviews on the state-of-the-art MARL algorithms in the context of CAVs. Therefore, this paper delivers a comprehensive review of the application of MARL techniques within the field of CAV control. The paper begins by introducing MARL, followed by a detailed explanation of its unique advantages in addressing complex mobility and traffic scenarios that involve multiple agents. It then presents a comprehensive survey of MARL applications on the extent of control dimensions for CAVs, covering critical and typical scenarios such as platooning control, lane-changing, and unsignalized intersections. In addition, the paper provides a comprehensive review of the prominent simulation platforms used to create reliable environments for training in MARL. Lastly, the paper examines the current challenges associated with deploying MARL within CAV control and outlines potential solutions that can effectively overcome these issues. Through this review, the study highlights the tremendous potential of MARL to enhance the performance and collaboration of CAV control in terms of safety, travel efficiency, and economy.
@@
Connected and automated vehicles, Multi-agent reinforcement learning, Intelligent transportation systems, Vehicle control.
@@ I Introduction
The transportation industry is currently experiencing a significant revolution, driven by the utilization of advanced techniques such as artificial intelligence (AI), the Internet of Things (IoT), and sensor technologies [1].
Notably, automated vehicles (AVs) represent a pivotal development in this ongoing transportation revolution with their potential to alleviate traffic congestion, improve road safety, and enhance accessibility [2]. Recently, the advent of connected and automated vehicles (CAVs) has taken this revolution a step further, combining the capabilities of AVs to establish a networked transportation ecosystem by using advanced communication technologies, such as vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and vehicle-to-cloud (V2C) [3, 4]. This interconnection and collaboration significantly augment the capabilities of individual AVs by extending their perception beyond immediate line-of-sight constraints and enhancing their collaborative decision-making abilities, thereby improving safety and traffic efficiency [5, 6, 7, 8]. However, the development of a robust CAV paradigm that tightly integrates computing, communication, and control technologies presents a considerable challenge, especially in the presence of complex interactions among AVs and intricate traffic scenarios involving other road users (e.g., bicyclists, pedestrians, and human-driven vehicles (HDVs)).
I-A Development of CAV control
@@
Control of CAVs stands as a key process within the above revolution and enables data-driven enhancements to transportation safety and sustainability by integrating seamlessly with diverse disciplines such as energy management, urban planning, and the broader social context [9, 10, 11]. Conventional optimization control methods have been proposed to tackle the CAV control problems [12, 13]. Specifically, an idealized model predictive control (MPC) approach is introduced in [14] to coordinate the behaviors of all following CAVs in a platoon, resulting in minimal control delay and efficient management of car-following behaviors to mitigate traffic oscillations. In [13], the coordination and control of the CAVs at road intersections is formulated as a non-convex optimal control problem and solved through semidefinite relaxation with randomization. However, the effectiveness of these approaches relies heavily on accurate system modeling [15], which is often not available. Furthermore, these methods commonly require online optimization that demands substantial computational resources that render these approaches impractical for real-time applications of CAV control [16].
@@
On the other hand, reinforcement learning (RL), a data-driven approach, has gained substantial attention within the research community due to its outstanding abilities in addressing sequential decision-making problems, such as gaming [17], robotics [18], behavioral planning [19, 20, 21, 22], intelligent energy management [23, 24]. Similarly, the AV control has begun exploring the potential of employing RL for various traffic scenarios [25, 26, 27]. For instance, in [26], a safe RL framework based on an enhanced double deep Q-Network (DDQN) [28] is proposed for highway lane-changing, resulting in zero collisions. In [25], an RL-based car-following model is introduced to optimize driving behavior, enhancing travel efficiency, fuel economy, and safety at signalized intersections in real-time performance. This approach featuring an effective reward function, demonstrates strong performance under varying traffic demands and traffic light cycles with different durations. Taking into account the role of HDVs in the control of AVs, Shi et al. employ a distributed proximal policy optimization (DPPO) control strategy that allows them to learn from and respond to the behavior of the HDVs, optimizing performance at both the local subsystem level and the broader mixed traffic context [29]. Qu et al. introduce a control approach utilizing the deep deterministic policy gradient (DDPG) algorithm aimed at reducing traffic fluctuations and enhancing fuel economy. However, their approaches only consider a single agent (i.e., AV) and exhibit poor generalization when multiple agents are involved. Additionally, their study focuses solely on individual vehicle control and does not leverage shared information [30].
I-B Multi-agent reinforcement learning for CAV control
In scenarios involving multiple agents, multi-agent reinforcement learning (MARL), a natural extension of single-agent RL, has emerged as a promising tool for optimizing the collective behaviors of multiple agents, showing significant potential for advancing the control strategies employed in the CAV ecosystem [31, 32, 33, 34, 35]. Fig. 1 illustrates a comprehensive multi-agent system, i.e., CAVs, and depicts a holistic approach to CAV operation, from sensory input and data processing to inter-agent communication and eventual decision-making and motion execution. On the left, the primary input sources include a variety of sensors such as GPS/INS, Camera, Radar, and Lidar. These sensors play a vital role in gathering real-time environmental data. The data is then processed through sensor fusion, which integrates multiple data sources to produce a consistent, accurate, and useful description of the environment. This processed data is crucial for tasks like localization (determining the vehicle’s precise position), perception (recognizing objects and obstacles), and environment prediction (forecasting possible changes in the surroundings). Furthermore, the system is equipped to use historical data and maps, ensuring that the vehicle’s decision-making is informed not only by real-time data but also by past experiences and pre-existing geographic information.
Figure 1: The development of connected and autonomous vehicles
The system’s architecture is anchored by multiple AVs, each endowed with its own repository for private data as well as access to a pool of shared information. These AVs are equipped with the capability to engage in dialogue and data exchange with neighboring units via local communication links, empowering them to enhance their decision-making processes and broaden their situational awareness. Next to this, the decision-making and motion planning module delineates the system’s logical and strategic framework. Based on the rich data inputs and dynamic interplay of inter-agent communications, decisions are formulated and motion trajectories are charted. The steering control and powertrain control modules demonstrate how these decisions translate into physical actions, such as turning the steering wheel or adjusting the engine’s or motor operation. A comprehensive examination and discourse on these components will be presented in Sections 3 and 4.
@@ I-B Multi-agent reinforcement learning for CAV control
Take some instances for example, a fully scalable and decentralized MARL algorithm is developed in [32] for large-scale traffic signal control, demonstrating superior control performance over the other state-of-the-art decentralized MARL algorithms. Furthermore, in [35], a scalable and safe MARL algorithm based on the parameter-sharing scheme is proposed for on-ramp merging in mixed traffic. Despite recent progress, the development of efficient and scalable MARL algorithms still presents significant challenges. The primary concern is guaranteeing the robustness and safety of these algorithms, especially in the face of diverse and uncertain traffic scenarios [35]. Furthermore, the dynamic behaviors of CAVs lead to non-stationary environments, adding layers of complexity to the development of MARL algorithms [32]. Another challenge lies in the lack of realistic simulators capable of accurately modeling traffic scenarios and other road users, which are essential for effective training of MARL algorithms [34]. Lastly, the real-world implementations of these algorithms remain an uphill task due to sim-to-real gaps [36, 35], and technological, legislative, and societal barriers [37]. While a few articles describe the potential of MARL in CAVs [38, 39, 40], none of them, to the best of our knowledge, are specifically devoted to the application of MARL for CAV control in intelligent transport systems (ITS) domain.
I-C Contributions of this review
@@
This paper is thus intended to provide a comprehensive and systematic review of MARL in the context of CAVs. We survey the recent developments in MARL algorithms, discussing their diverse applications in aspects of CAV control based on the extent of control dimensions, such as platooning control, lane changing, on-ramp merging, and traffic signal cooperation. We also aim to highlight and delve into the technical challenges that MARL faces in the context of these applications and discuss potential research directions that could be pursued to address these challenges. This review not only serves to provide a clear, up-to-date understanding of the current state-of-the-art MARL algorithms applied in CAV control, but it also outlines the potential direction for future work in this area. The main contributions of this paper are summarized as follows.
@@
    A self-contained overview of RL and MARL background, specially tailored for the automotive community, is provided, recognizing that these concepts may not be widely familiar within this specific domain.
@@
    We review recent advances in MARL for CAV control, proposing a new taxonomy of MARL algorithms based on the degree of control signals. Furthermore, we provide an extensive examination of the leading simulation platforms employed in MARL research for CAVs.
@@
    We discuss the remaining challenges and potential future directions, shedding light on the exciting opportunities that lie ahead in this field.
@@
    We establish a well-classified list of relevant papers, which will be actively updated and can be accessed at the following site: https://github.com/huahuaedi/MARL_in_CAV_control_review.
The rest of this paper is organized as follows: Sec. II provides the preliminaries of RL and MARL; a detailed review of MARL applied in CAV control from different control dimensions and mainstream simulation platforms is described in Sec. III. Sec. IV summarizes the remaining challenges and opportunities. Finally, Sec. V concludes the review.
II Backgrounds
In this section, we offer an in-depth overview of the core principles of RL. Subsequently, we explore several leading MARL algorithms, establishing the context to enhance the understanding of our review.
II-A Preliminaries of reinforcement learning (RL)
RL, often mathematically formulated as a Markov Decision Process (MDP), has emerged as a promising data-driven method for sequential decision-making [41]. Recent advancements in Deep Neural Networks (DNNs) have further bolstered RL’s capabilities in handling intricate tasks. Noteworthy examples of these algorithms include the deep Q-network (DQN) [42], deep deterministic policy gradient (DDPG) [43], and advantage actor-critic (A2C) [44]. For instance, the AlphaGo computer program, based on DQN, made history by becoming the first of its kind to defeat a professional human player in Go and subsequently beat a world champion in the game [42]. Additionally, in a study by Lee et al. [18], researchers successfully applied Trust Region Policy Optimization (TRPO, [45]) to navigate quadrupedal robots across challenging terrains, including surfaces like mud and snow, and through dynamic footholds.
In an RL framework (refer to Figure 2), the learner (referred to as the agent) traverses the environment employing a trial-and-error methodology. The agent makes decisions, executes corresponding actions within the environment, and, in turn, receives a reward signal along with a new state. This reward, conveyed by the environment, offers feedback to the agent, indicating whether the impact of its actions is positive or negative. Mathematically, the RL problem can be formulated as a Markov Decision Process (MDP). The MDP $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$ is defined as follows:
  1. 1.
    State space $\mathcal{S}$: A comprehensive set of states describing the environment, specifying the position and conditions of agents at a given time step $t$. An observation $o$ provides a partial state description and may lack complete information. The environment is fully observed if the agent can access the entire state; otherwise, it is partially observed.
  2. 2.
    Action space $\mathcal{A}$: A collection of valid actions within the environment. Environments, such as Atari and Go, have discrete action spaces, limiting the agent to a finite set of moves. Environments controlling a physical robot have continuous action spaces, where actions are real-valued vectors.
  3. 3.
    Reward $\mathcal{R}(s_{t},a_{t},s_{t+1})$:The environment returns a reward upon executing action $a_{t}$ at state $s_{t}$. The reward signal’s value may be positive or negative, depending on the agent’s actions.
  4. 4.
    Transition Probability $\mathcal{P}_{ss^{\prime}}(S_{t+1}=s^{\prime}|S_{t}=s)$: the transition probability describes the likelihood of an agent moving from one state to another.
@@ II-A Preliminaries of reinforcement learning (RL)
In the subsequent subsections, we will explore three prevalent RL algorithms: Deep Q-learning, Policy Gradient, and the Actor-Critic Network.
II-A1 Deep Q-Learning
In Q-learning, the Q-function, denoted as $Q_{\theta}$, is often parameterized by a set of parameters $\theta$. This parameterization can be achieved using various function approximators, such as Q-tables [46], Linear Regression (LR) [47], and Deep Neural Networks (DNNs) [42]. The temporal difference, denoted as $(\mathcal{T}Q_{\theta^{-}}-Q_{\theta})(s_{t},a_{t})$, forms the basis for updating $\theta$. Here $\mathcal{T}$ and $Q_{\theta^{-}}$ represent the dynamic programming (DP) operator and a recently frozen model $\theta^{-}$ [4], respectively. To reduce variance in estimating Q-values and improve exploration, techniques like the $\epsilon-\text{greedy}$ and experience replay are commonly integrated into deep Q-learning [47]. We can obtain the optimal action $a^{*}(s)$ through the $Q^{*}(s_{t}=s,a_{t}=a)$ as:

### v3

@@ Abstract
Connected and automated vehicles (CAVs) are considered as a potential solution for future transportation challenges, aiming to develop systems that are efficient, safe, and environmentally friendly. However, CAV control presents significant challenges due to the complexity of interconnectivity and coordination required among vehicles. Multi-agent reinforcement learning (MARL), which has shown notable advancements in addressing complex problems in autonomous driving, robotics, and human-vehicle interaction, emerges as a promising tool to enhance CAV capabilities. Despite its potential, there is a notable absence of current reviews on mainstream MARL algorithms for CAVs. To fill this gap, this paper offers a comprehensive review of MARL’s application in CAV control. The paper begins with an introduction to MARL, explaining its unique advantages in handling complex and multi-agent scenarios. It then presents a detailed survey of MARL applications across various control dimensions for CAVs, including critical scenarios such as platooning control, lane-changing, and unsignalized intersections. Additionally, the paper reviews prominent simulation platforms essential for developing and testing MARL algorithms. Lastly, it examines the current challenges in deploying MARL for CAV control, including macro-micro optimization, communication, mixed traffic, and sim-to-real challenges. Potential solutions discussed include hierarchical MARL, decentralized MARL, adaptive interactions, and offline MARL.
Note to Practitioners—This paper presents multi-agent reinforcement learning as a solution to the challenges of controlling connected and automated vehicles in complex scenarios like platooning, lane changes, and intersections. MARL offers a more adaptive approach than traditional methods, potentially improving traffic flow, safety, and fuel efficiency in real-world settings and leading to more efficient and reliable transportation systems. The paper also reviews current MARL algorithms and simulation platforms, providing a resource for implementing these advanced control strategies in practice. However, the deployment of MARL in real-world CAV systems is still in its early stages and faces challenges such as ensuring reliable communication between vehicles and managing mixed traffic environments that include both automated and human-driven vehicles. Future research is needed to address these challenges and validate the approach in diverse and dynamic traffic conditions. This paper serves as a stepping stone for practitioners aiming to develop more reliable and adaptive CAV systems in the near future.
@@
Connected and automated vehicles, multi-agent reinforcement learning, intelligent transportation systems, and Vehicle control.
@@ I Introduction
The transportation industry is experiencing a transformative shift driven by advancements in automation, artificial intelligence (AI), the Internet of Things (IoT), and sensor technologies [1, 2]. A central element of this shift is the development of automated vehicles (AVs), which promise to mitigate traffic congestion, enhance road safety, and increase accessibility [3, 4]. The evolution of this field has led to the emergence of connected and automated vehicles (CAVs), which integrate AVs into a cohesive, networked transportation system through advanced communication technologies such as vehicle-to-vehicle (V2V), vehicle-to-infrastructure (V2I), and vehicle-to-cloud (V2C) [5]. This integration significantly expands the operational capabilities of AVs, enabling extended perception, enhanced collaborative decision-making, and improved traffic efficiency [6, 7]. However, the realization of an effective CAV framework that harmonizes control, computing, and communication presents substantial challenges, particularly in managing the intricate interactions between AVs and other road users, including pedestrians, cyclists, and human-driven vehicles (HDVs) [8]. Addressing these challenges is critical for advancing automation in transportation and ensuring the safe, efficient deployment of CAV systems.
@@
Control of CAVs is pivotal in enhancing transportation safety and sustainability, integrating with disciplines like energy management, urban planning, and social contexts [9, 10]. Conventional optimization methods have addressed CAV control challenges, such as using model predictive control (MPC) to coordinate platoon behavior, minimizing control delay, and reducing traffic oscillations [11]. Additionally, non-convex optimal control problems, such as CAV coordination at intersections, have been tackled with semidefinite relaxation techniques [12]. However, these approaches often depend on precise system modeling, which is not always available [13], and require substantial computational resources, making them impractical for real-time CAV control [14].
@@
Besides, reinforcement learning (RL) has gained increasing attention within the research field due to its outstanding abilities in addressing sequential decision-making tasks, such as gaming [15], robotics [16, 17], behavioral planning [18], intelligent energy management [19, 20, 21]. Similarly, the AV control has begun exploring the potential of employing RL for various traffic scenarios [22, 23]. For instance, in [22], a safe RL framework based on an improved double deep Q-Network (DDQN) [24] is proposed for highway lane-changing, resulting in zero collisions. In [25], a RL-based model for plattoning control is introduced to enhance fuel economy, driving efficiency, and safety at signalized intersections through real-time optimization. This approach featuring an effective reward function, demonstrates strong performance under varying traffic demands and traffic light cycles with different durations. Taking into account the role of HDVs in the control of AVs, Shi et al. employ a distributed proximal policy optimization (DPPO) control strategy that allows them to learn from and respond to the behavior of the HDVs, optimizing performance at both the local subsystem level and the broader mixed traffic context [26]. Qu et al. introduce a control approach utilizing the deep deterministic policy gradient (DDPG) algorithm aimed at reducing traffic fluctuations and enhancing fuel economy. However, these approaches only consider a single agent (i.e., AV) and exhibit poor generalization when multiple agents are involved. Additionally, their study focuses solely on individual vehicle control and does not leverage shared information [27].
Multi-agent reinforcement learning (MARL) represents a significant research direction within artificial intelligence, extending the concepts of single-agent RL to scenarios involving multiple interacting agents. This innovative approach has found applications across a broad spectrum of fields. Notably, the gaming and simulation industry has utilized MARL to create more complex and interactive environments [28]. Furthermore, within the realm of finance, MARL models are employed to simulate the dynamics of markets and the behaviors of agents [29]. In traffic control and management, MARL algorithms have been instrumental in optimizing traffic flows and reducing congestion. The realm of AVs and robotics has also benefited greatly, with MARL enabling cooperative navigation and coordination among vehicles and facilitating collaborative tasks in AVs and robotics [30]. Extended to CAV control, each vehicle acts as an independent agent that learns from its interactions with the surrounding environment, which includes other vehicles, traffic signals, and road conditions. Each CAV learns not only to maneuver safely and efficiently on its own but also to cooperate and communicate with other CAVs [31]. Therefore, MARL offers a valuable strategy for improving the performance and efficiency of CAVs, paving the way for more sophisticated, collaborative, and flexible transportation networks.
In scenarios involving multiple agents, MARL has emerged as a promising tool for optimizing the collective behaviors, showing significant potential for advancing the control strategies employed in the CAV ecosystem [32]. Fig. 1 illustrates a comprehensive multi-agent system, i.e., CAVs, and depicts a holistic approach to CAV operation, from sensory input and data processing to inter-agent communication and eventual decision-making and motion execution. On the left, real-time environmental data from various sensors, including GPS/INS, cameras, radar, and lidar, is gathered and then integrated via sensor fusion. Within multi-agent system, each AV maintains its own data repository while accessing shared information, facilitating inter-agent communication for enhanced decision-making and situational awareness [33]. This system includes modules for decision-making and motion planning, which use data and inter-vehicle communication to make decisions and plan movements. Finally, the steering control and powertrain control modules demonstrate how these decisions translate into physical actions, such as turning the steering wheel or adjusting the engine’s or motor operation.
Figure 1: The multi-agent control system for CAVs: the left side represents the information inputs of the overall control system, while the right side depicts the studied multi-agent control system for CAVs.
@@
A good example of large-scale traffic signal control is developed [34] through a highly scalable and decentralized MARL algorithm, demonstrating superior control performance over the other leading decentralized MARL algorithms. Furthermore, in [32], the challenge of on-ramp merging in mixed traffic scenarios has been explored with a scalable and safe MARL algorithm that utilizes a parameter-sharing technique to enhance safety and scalability. Despite recent progress, the development of efficient and scalable MARL algorithms still presents significant challenges. The primary concern is guaranteeing the robustness and safety of these algorithms, especially in the face of diverse and uncertain traffic scenarios [32]. Furthermore, the dynamic behaviors of CAVs lead to non-stationary environments, adding layers of complexity to the development of MARL algorithms [34]. Another challenge lies in the lack of realistic simulators capable of accurately modeling traffic scenarios and other road users, which are essential for effective training of MARL algorithms [35]. Lastly, the real-world implementations of these algorithms remain an uphill task due to sim-to-real gaps [32], and technological, legislative, and societal barriers . While a few articles describe the potential of MARL in CAVs [36, 37], none of them, to the best of our knowledge, are specifically devoted to the application of MARL for CAV control.
@@
This paper is thus intended to deliver a comprehensive and systematic review of MARL within the realm of CAVs, such as lane changing, platooning control, traffic signal cooperation, and on-ramp merging. Through this review, we not only provide a clear, up-to-date understanding of the current classic MARL algorithms applied in CAV control, but we also outline the potential direction for future work in this area. The main contributions of the work presented in this paper are summarized as follows:
@@
    We survey the recent developments in MARL algorithms, discussing their diverse applications in aspects of CAV control based on the extent of control dimensions. Furthermore, we provide an extensive examination of the leading simulation platforms employed in MARL research for CAVs.
@@
    We highlight and explore the technical challenges that MARL faces in these applications and discuss potential research directions to address these challenges.
@@
    The work, including a well-classified list of relevant papers, has been presented at the following site: https://github.com/huahuaedi/MARL_in_CAV_control_review.
This paper is organized as follows: Section II introduces the basics of RL and MARL; A detailed review of MARL applied in CAV control from different control dimensions and mainstream simulation platforms are described in section III. Section IV summarizes the remaining challenges and opportunities. Finally, Section V concludes the review.
II Methodologies
In this section, we begin with a comprehensive overview of the fundamental principles of RL. We then delve into various prominent MARL algorithms, establishing the context to enhance the understanding of our review.
II-A Preliminaries of RL
RL, frequently modeled through a mathematical framework known as a Markov Decision Process (MDP), has risen as an effective method for data-driven sequential decision-making [38]. Recently, deep neural networks (DNN) have substantially elevated the capability of RL to manage complex problems [39]. Key developments include sophisticated algorithms such as the Deep Q-Network (DQN) [39], Deep Deterministic Policy Gradient (DDPG) [40], and Advantage Actor-Critic (A2C) [41]. For instance, the AlphaStar, which operates on principles similar to DQN, marked a significant milestone by outperforming professional esports players in StarCraft II, illustrating the power of RL in strategic game environments [42]. Moreover, the development of the autonomous driving system by leveraging DDPG exemplifies the potential of RL, showcasing how vehicles can autonomously navigate through environments with dynamic obstacles [43].
In the context of a RL framework (as shown in Fig. 2), a learner known as the agent navigates through the environment using a process of trial and error. This agent assesses situations, undertakes specific actions in the environment, and consequently receives feedback in the form of a reward signal and a new state. The feedback, provided by the environment, serves as an indicator to the agent about the effectiveness of its actions, signaling whether they have a positive or negative impact. Many RL problems are framed as MDP, which provide a mathematical framework for modeling decision making where outcomes are partly random and partly under the control of a decision maker. MDPs are characterized by states, actions, transition dynamics, and rewards with $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$, where it is defined as follows:
  1. 1.
[…]
  2. 2.
    Action space $\mathcal{A}$: the set of possible moves or decisions the agent can make in a given state. The set of actions available can depend on the state. The action spaces fall into two categories: discrete or continuous.
  3. 3.
    Reward $\mathcal{R}(s_{t},a_{t},s_{t+1})$: a scalar feedback signal given to the agent from the environment after performing action $a_{t}$ in state $s_{t}$. The agent’s objective is to maximize the cumulative reward over time.
@@ II-A Preliminaries of RL
    Transition Probability $\mathcal{P}_{ss^{\prime}}(S_{t+1}=s^{\prime}|S_{t}=s)$: the chance of an agent transitioning from one state to a different state.
@@
In the following subsections, we will explore three prevalent RL algorithms, including Value-Based Methods, Policy-Based Methods, and Actor-Critic Methods.
II-A1 Value-Based Methods
In the domain of value-based methods within RL, the Q-function, represented as $Q_{\theta}$, is usually parameterized by a set of parameters denoted as $\theta$. This parameterization facilitates the use of different function approximators for estimating Q-values, including traditional Q-tables [44] and the complex Deep Neural Networks (DNN) [39]. The core mechanism for updating these parameters relies on the temporal difference (TD), expressed as $(\mathcal{T}Q_{\theta^{-}}-Q_{\theta})(s_{t},a_{t})$, where $\mathcal{T}$ signifies the dynamic programming operator and $Q_{\theta^{-}}$ represents a version of the model parameters $\theta^{-}$ that is temporarily fixed to stabilize updates [45]. To mitigate the variance in Q-value estimates and enhance the exploration ability of the algorithm, strategies such as the $\epsilon-\text{greedy}$ method and experience replay mechanism are frequently employed within deep Q-learning frameworks [46]. The optimal action $a^{*}(s)$ can be described as: